# DEXTER: Diffusion-Guided EXplanations with TExtual Reasoning for Vision Models

### v1

@@ 4.2.3 Bias explanation
To evaluate DEXTER’s ability to identify and explain biases in classifiers, we conduct an analysis using the FairFaces dataset [14]. Specifically, we train two binary classifiers to distinguish between two age groups: 20-29 (class 1) and 50-59 (class 2). These classifiers are trained on two variants of the FairFaces dataset: (1) a balanced dataset with equal male and female representation in both classes, (2) a dataset where males are overrepresented in class 1, and females are overrepresented in class 2. These variations allowed us to systematically assess how biases in the training data are reflected in the classifier’s behavior and how well DEXTER captures these biases.
@@
1. User Study. In addition to the earlier study, participants evaluated and ranked DEXTER’s bias reports for two classifiers. Informed of dataset biases, they assessed report accuracy, clarity, and interpretability. We report human Mean Opinion Scores (MOShumans, 1–5 scale), and complement them with automated evaluations: MOSLLM from a VLM and G-eval [20] metrics, including G-evalcon for consistency with classifier behavior. These scores jointly assess the linguistic, structural, and explanatory quality of DEXTER’s outputs.
2. Comparison with training-derived reports. We compare DEXTER’s data-free reports to those generated from training set images using the same pipeline (captioning + LLM-based reporting). Their similarity is quantified via sentence transformer similarity (STS) [34], measuring alignment with data-grounded biases.
@@ D.1 Quantitative evaluation metrics: CLIP-IQA and Semantic CLIP-IQA
CLIP-IQA [45] is a metric originally introduced to evaluate image quality. In this work, we leverage CLIP-IQA to compare the quality of images generated by DEXTER against those produced by other approaches. Specifically, CLIP-IQA calculates the similarity between generated images and two fixed prompts, returning the probability that an image is closer in similarity to the first prompt rather than the second. The fixed prompts used in the original formulation are “Good photo.” and “Bad photo.
@@
Yosinski et al. [47] 0.37 $\pm$ 0.19 0.33 $\pm$ 0.23
Mahendran and Vedaldi [23] 0.82 $\pm$ 0.15 0.72 $\pm$ 0.15
Nguyen et al. [26] 0.74 $\pm$ 0.21 0.57 $\pm$ 0.30
DiffExplainer [31] 0.89 $\pm$ 0.09 0.90 $\pm$ 0.05
@@ Appendix E Slice Discovery Details
DEXTER identifies the words explaining an input visual classifier, as presented in Sect. 3. Specifically, given a classifier $f$ and a target class $c$, it optimizes a text prompt to extract a set of top-k class words $W_{C}=w_{c1},...,w_{ck}$ (Tab. 8). Each word $w_{ci}$ is encoded using CLIP’s text encoder to obtain an embedding $t_{ci}$ and the class prototype is defined as the average embedding: $\bar{t_{c}}=\frac{1}{k}\sum t_{ci}$ Then, following the approach in B2T [16], at inference time, each image $x_{j}$ of the dataset is encoded into $v_{j}$ using CLIP’s image encoder. The similarity between the image and the class is then computed as:
@@
If an image has an higher similarity with the class-words that correspond to its real class it is labeled as unbiased slice (e.g. an image belonging to class landbirds with an higher similarity with landbirds class-words), while if an image has an higher similarity with its counterpart class it is labeled as biased slice (e.g. an image belonging to class landbirds has an higher similarity with waterbirds class-words). Fig. 7 presents the ROC curves obtained using the CLIP-based matching approach on the Waterbirds dataset, in comparison with other SOTA slice discovery methods [16, 8, 12]. The results illustrate how effectively the training images are partitioned into the four groups (waterbird on water, waterbird on land, landbird on land, landbird on water) by the DEXTER-derived words. Finally, we follow the debiased training scheme from [36] to train a debiased classifier on the dataset.
In all our slice discovery experiments, we set $k=4$ both for CelebA and Waterbirds. We used 1000 DEXTER optimization steps and single word prompting (see Sect. 4.2.4) as an activation maximization strategy to discover the above words. Tab. 8 reports the discovered words for both datasets, where class 0 of CelebA has been left blank as in B2T [16].
@@ F.2 Bias Report Quality Assessment Metrics
In this section, we provide additional explanations for the metrics used to evaluate the quality of DEXTER’s bias reports, as shown in Table 2. Following Liu et al. [20], we use the G-eval consistency metric, which involves prompting a large language model (LLM) to assess a text’s factualness, scoring each report on a scale from 1 to 5. Since the original prompts for these metrics were designed for summarization, we create the following G-eval Consistency Prompt to better align with our task.
@@
For both the G-eval and $MOS_{\text{LLM}}$ configurations, we adopt the same settings as Liu et al. [20]. Specifically, we use gpt-4-0613 with a temperature of 2, max_tokens = 0, top_p = 1.0, frequency_penalty = 0, and presence_penalty = 0.
To enhance response diversity, we generate $n=20$ outputs. Furthermore, in both G-eval and MOS LLM system prompts, we provide the same evaluation instructions that were given to human evaluators during the user study, also following the best practices to reduce hallucinations provided by [4].
@@ 59. Justification:
    Asset Type License Github / URL Citation/Reference SalientImagenet data non-commercial research singlasahil14/salient_imagenet [41] Waterbirds data non-commercial research kohpangwei/group_DRO [36] CelebA data non-commercial research https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html [21] FairFaces data CC BY 4.0 joojs/fairface [14] BERT Pretrained Model Apache license 2.0 https://huggingface.co/google-bert/bert-base-uncased [5] CLIP Pretrained Model MIT license openai/CLIP [32] Stable Diffusion 1.5 Pretrained Model CreativeML Open RAIL-M https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 [35] GPT-4o LLM API-based use under OpenAI terms https://platform.openai.com/docs/overview OpenAI

### v2

@@ 4.2.3 Bias explanation
To evaluate DEXTER’s ability to identify and explain biases in classifiers, we conduct an analysis using the FairFaces dataset [karkkainen2021fairface]. Specifically, we train two binary classifiers to distinguish between two age groups: 20-29 (class 1) and 50-59 (class 2). These classifiers are trained on two variants of the FairFaces dataset: (1) a balanced dataset with equal male and female representation in both classes, (2) a dataset where males are overrepresented in class 1, and females are overrepresented in class 2. These variations allowed us to systematically assess how biases in the training data are reflected in the classifier’s behavior and how well DEXTER captures these biases.
@@
1. User Study. In addition to the earlier study, participants evaluated and ranked DEXTER’s bias reports for two classifiers. Informed of dataset biases, they assessed report accuracy, clarity, and interpretability. We report human Mean Opinion Scores (MOShumans, 1–5 scale), and complement them with automated evaluations: MOSLLM from a VLM and G-eval [liu2023g] metrics, including G-evalcon for consistency with classifier behavior. These scores jointly assess the linguistic, structural, and explanatory quality of DEXTER’s outputs.
2. Comparison with training-derived reports. We compare DEXTER’s data-free reports to those generated from training set images using the same pipeline (captioning + LLM-based reporting). Their similarity is quantified via sentence transformer similarity (STS) [reimers2019sentence], measuring alignment with data-grounded biases.
@@ D.1 Quantitative evaluation metrics: CLIP-IQA and Semantic CLIP-IQA
CLIP-IQA [wang2022exploring] is a metric originally introduced to evaluate image quality. In this work, we leverage CLIP-IQA to compare the quality of images generated by DEXTER against those produced by other approaches. Specifically, CLIP-IQA calculates the similarity between generated images and two fixed prompts, returning the probability that an image is closer in similarity to the first prompt rather than the second. The fixed prompts used in the original formulation are “Good photo.” and “Bad photo.
@@
yosinski2015understanding 0.37 $\pm$ 0.19 0.33 $\pm$ 0.23
mahendran2016visualizing 0.82 $\pm$ 0.15 0.72 $\pm$ 0.15
nguyen2016synthesizing 0.74 $\pm$ 0.21 0.57 $\pm$ 0.30
DiffExplainer [pennisi2024diffexplainer] 0.89 $\pm$ 0.09 0.90 $\pm$ 0.05
@@ Appendix E Slice Discovery Details
DEXTER identifies the words explaining an input visual classifier, as presented in Sect. 3. Specifically, given a classifier $f$ and a target class $c$, it optimizes a text prompt to extract a set of top-k class words $W_{C}=w_{c1},...,w_{ck}$ (Tab. 8). Each word $w_{ci}$ is encoded using CLIP’s text encoder to obtain an embedding $t_{ci}$ and the class prototype is defined as the average embedding: $\bar{t_{c}}=\frac{1}{k}\sum t_{ci}$ Then, following the approach in B2T [kim2024discovering], at inference time, each image $x_{j}$ of the dataset is encoded into $v_{j}$ using CLIP’s image encoder. The similarity between the image and the class is then computed as:
@@
If an image has an higher similarity with the class-words that correspond to its real class it is labeled as unbiased slice (e.g. an image belonging to class landbirds with an higher similarity with landbirds class-words), while if an image has an higher similarity with its counterpart class it is labeled as biased slice (e.g. an image belonging to class landbirds has an higher similarity with waterbirds class-words). Fig. 7 presents the ROC curves obtained using the CLIP-based matching approach on the Waterbirds dataset, in comparison with other SOTA slice discovery methods [kim2024discovering, eyubogludomino, jaindistilling]. The results illustrate how effectively the training images are partitioned into the four groups (waterbird on water, waterbird on land, landbird on land, landbird on water) by the DEXTER-derived words. Finally, we follow the debiased training scheme from [sagawa2019distributionally] to train a debiased classifier on the dataset.
In all our slice discovery experiments, we set $k=4$ both for CelebA and Waterbirds. We used 1000 DEXTER optimization steps and single word prompting (see Sect. 4.2.4) as an activation maximization strategy to discover the above words. Tab. 8 reports the discovered words for both datasets, where class 0 of CelebA has been left blank as in B2T [kim2024discovering].
@@ F.2 Bias Report Quality Assessment Metrics
In this section, we provide additional explanations for the metrics used to evaluate the quality of DEXTER’s bias reports, as shown in Table 2. Following liu2023g, we use the G-eval consistency metric, which involves prompting a large language model (LLM) to assess a text’s factualness, scoring each report on a scale from 1 to 5. Since the original prompts for these metrics were designed for summarization, we create the following G-eval Consistency Prompt to better align with our task.
@@
For both the G-eval and $MOS_{\text{LLM}}$ configurations, we adopt the same settings as liu2023g. Specifically, we use gpt-4-0613 with a temperature of 2, max_tokens = 0, top_p = 1.0, frequency_penalty = 0, and presence_penalty = 0.
To enhance response diversity, we generate $n=20$ outputs. Furthermore, in both G-eval and MOS LLM system prompts, we provide the same evaluation instructions that were given to human evaluators during the user study, also following the best practices to reduce hallucinations provided by [understanding_bias].
@@ 59. Justification:
    Asset Type License Github / URL Citation/Reference SalientImagenet data non-commercial research singlasahil14/salient_imagenet [singla2022salient] Waterbirds data non-commercial research kohpangwei/group_DRO [sagawa2019distributionally] CelebA data non-commercial research https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html [liu2015faceattributes] FairFaces data CC BY 4.0 joojs/fairface [karkkainen2021fairface] BERT Pretrained Model Apache license 2.0 https://huggingface.co/google-bert/bert-base-uncased [devlin2018bert] CLIP Pretrained Model MIT license openai/CLIP [radford2021learning] Stable Diffusion 1.5 Pretrained Model CreativeML Open RAIL-M https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 [rombach2022high] GPT-4o LLM API-based use under OpenAI terms https://platform.openai.com/docs/overview OpenAI